Could I please get some help with this?
1/2(2x+2a)=1/3(21+3b) For x

Answer by stanbon(48568)

1/2(2x+2a)=1/3(21+3b) For x
---------
x+a = 7+b
x=7+b-a
=========
Cheers,
Stan H.